An American flag has a perimeter of 28ft. Find the dimensions of the flag if the length is 4 ft less than twice its width.
Answer by mananth(16946)

let width be x
length = 2x-4
..
Perimeter = 2*(L+W)
28=2(x+2x-4)
28=6x-8
add 8
28+8=6x+8-8
36=6x
divide by 6
36/6= 6x/6
6=x
width = 6 feet
length = 2x-4
=2*6-4
=12-4
=8 feet